# Color-avoiding Ramsey for directed paths and vector sequences

### v1

@@ 3. Simple proofs
The first two authors [20, Lemma 1.3] proved the following quantitative variant of the triangle removal lemma in tournaments: There is a constant $c>0$ such that if a tournament on $N$ vertices is $\delta$-far from transitive, then it contains at least $c\delta^{2}N^{3}$ cyclic triangles. For each cyclic triangle, arbitrarily label the edges cyclically as $1,2,3$. Since there are $q$ colors, there are $q^{3}$ possible colorings of the edges of a cyclic triangle with the edges labeled as $1,2,3$. By the pigeonhole principle, there are at least $c\delta^{2}N^{3}/q^{3}$ cyclic triangles whose labeled edges have the same color pattern (call this pattern $P$). Greedily delete edges from the tournament that are in fewer than $2c\delta^{2}N/q^{3}$ such cyclic triangles of color pattern $P$. As there are fewer than $N^{2}/2$ edges, we still have more than $c\delta^{2}N^{3}/q^{3}-(N^{2}/2)(2c\delta^{2}N/q^{3})>0$ cyclic triangles of color pattern $P$ remaining after we are done with this process. Each edge that remains (and there is at least one since there is at least one cyclic triangle remaining) extends to at least $2c\delta^{2}N/q^{3}$ cyclic triangles of color pattern $P$. Starting with the vertices $v_{1},v_{2},v_{3}$ of one remaining cyclic triangle of color pattern $P$ with $v_{1}\rightarrow v_{2}$, we can greedily build the desired directed path (and furthermore the square of a path with the desired properties) by adding, after $v_{i-1},v_{i}$, a vertex $v_{i+1}$ not already on the path such that $v_{i-1},v_{i},v_{i+1}$ forms a cyclic triangle of color pattern $P$ in what remains. We can necessarily find such a vertex $v_{i+1}$ as long as $2c\delta^{2}N/q^{3}>i-1$ since each edge on the path extends to more than $2c\delta^{2}N/q^{3}$ cyclic triangles of color pattern $P$ in what remains. Hence, the path length we get in the end is at least $2c\delta^{2}N/q^{3}$. ∎
@@ 3.2. General reductions and bounds
If $K$ is a $q$-edge-colored ordered complete graph, we denote by $f_{q,r}(K)$ the maximum length of a monotone path in $K$ which is colored by at most $r$ colors. We then let $f_{q,r}(N)$ be the minimum of $f_{q,r}(K)$ over all $q$-edge-colored ordered $N$-vertex complete graphs $K$. The following simple observation, which already appears in [22, 26] captures the fact that when r=q1r=q-1, this notion is equivalent to the study of (q1)(q-1)-increasing sequences in $n]^{q}$.
Proposition 3.1.
We have $f_{q,q-1}(N)\leqslant n$ if and only if $F_{q,q-1}(n)\geqslant N$.
Proof.
First, suppose that $f_{q,q-1}(N)\leqslant n$. Thus, there exists some $q$-edge-colored ordered clique $K$ on vertex set $[N]$ in which every monotone path with at most q1q-1 colors contains at most $n$ vertices. We define vectors $x_{1},\dots,x_{N}$ by letting $x_{a}$ record, for each $i\in[q]$, the length of the longest monotone path ending at $a$ which does not use color $i$. Note that we are measuring path length by the number of vertices, hence each entry of each vector $x_{a}$ is at least $1$. Moreover, by assumption, every color-avoiding monotone path in $K$ has length at most $n$, so we conclude that $x_{a}\in[n]^{q}$ for all $a\in[N]$. We now claim that for all $a<b$, we have that $x_{a}<_{q-1}x_{b}$. Indeed, if the color of $ab$ is $i$, then for every $j\neq i$, we can extend every $j$-avoiding monotone path ending at $x_{a}$ by one to be a $j$-avoiding monotone path ending at $x_{b}$. Hence, every coordinate of $x_{b}$, except potentially the $i$th coordinate, is strictly larger than the corresponding coordinate of $x_{a}$. We conclude that $F_{q,q-1}(n)\geqslant N$, as desired.
For the converse, let’s suppose that $F_{q,q-1}(n)\geqslant N$, and fix vectors $x_{1},\dots,x_{N}\in[n]^{q}$ witnessing this. We construct a $q$-edge-coloring of the complete graph on vertex set $[N]$ as follows: for every $a<b$, we know that $x_{a}<_{q-1}x_{b}$, hence there is at most one coordinate $i$ for which $(x_{a})_{i}\geqslant(x_{b})_{i}$. If there is such an $i$, we color the edge $ab$ with color $i$; if not, we color it arbitrarily. The key property of this coloring is that if $ab$ is not colored $i$, then $(x_{a})_{i}<(x_{b})_{i}$ (for if this were not the case, we would be forced to color the edge $i$ by the rule above). We claim that this coloring has no monotone color-avoiding path of length $n+1$, proving that $f_{q,q-1}(N)\leqslant n$. Indeed, suppose for contradiction that $v_{0},v_{1},\dots,v_{n}$ form a monotone color-avoiding path in this coloring, say avoiding the color $i$. Since each consecutive pair is not colored $i$, we must have that $(x_{v_{a}})_{i}<(x_{v_{a+1}})_{i}$ for all $0\leqslant a<n$. But the numbers $(x_{v_{0}})_{i},\dots,(x_{v_{n}})_{i}$ all lie in $[n]$, so it is not possible for these $n+1$ numbers to be strictly increasing. This contradiction completes the proof. ∎
It is worth remarking that the same argument above shows that $F_{q,r}(n)$ is equal to the largest $N$ for which the following holds. There is a complete ordered graph on $N$ vertices in which each edge is assigned a subset of $r$ colors from a set of $q$ colors, and no color appears in the color set of every edge of a monotone path $P$ of length $n+1$. This is one less than the set-coloring ordered Ramsey number of a monotone path with $n+1$ vertices. Set-coloring Ramsey numbers have been extensively studied recently for cliques (see [10, 12, 2]) and have found interesting connections to coding theory, and have also been studied for certain other classes of graphs (see e.g. [7] and the references therein). Ordered Ramsey numbers of graphs have also been extensively studied recently (see [11, 5] and the recent survey [4]).
Similarly, $G_{q,r}(n)$ is at most the largest $N$ for which the following holds. There is a tournament on $N$ vertices in which each edge is assigned a subset of $r$ colors from a set of $q$ colors, and no color appears in the color set of every edge of a directed path of length $n+1$. This is one less than the set-coloring oriented Ramsey number of a path with $n+1$ vertices with parameters $(q,r)$. Directed Ramsey numbers have similarly received much recent attention (see, for instance, [8, 18]). It would be interesting to study set-coloring Ramsey numbers for other graphs, ordered graphs, and acyclic digraphs.
Another useful observation is that monotone paths behave very cleanly under lexicographic product. Recall that an ordered graph is a graph equipped with a linear order on its vertex set. Given two ordered graphs $G_{1},G_{2}$, we define their lexicographic product $G_{1}\otimes G_{2}$ by first duplicating each vertex of $G_{2}$ to $\lvert G_{1}\rvert$ vertices, and replacing each edge of $G_{2}$ by a complete bipartite graph between these blowup sets. We then insert a copy of $G_{1}$ into each of these blowup sets, while otherwise maintaining the order. That is, each original vertex of $G_{2}$ now corresponds to an interval of length $\lvert G_{1}\rvert$ in the new ordering.
@@
One consequence of Proposition 3.3 is the upper bound $f_{q,r}(N)\leqslant N^{r/q}$ for all $N$ which are a power of $q$. Indeed, let $N=m^{q}$, and let $K_{1},\dots,K_{q}$ be ordered cliques on $m$ vertices, where all edges in $K_{i}$ are given color $i$. By Proposition 3.3, for every $S\subseteq[q]$, the length of the longest $S$-colored monotone path in $K=K_{1}\otimes\dots\otimes K_{q}$ is precisely $m^{\lvert S\rvert}$; in particular, the longest monotone path in $K$ receiving at most $r$ colors has length exactly $m^{r}=(m^{q})^{r/q}$, proving the bound $f_{q,r}(N)\leqslant N^{r/q}$ for $N=m^{q}$.
Another immediate corollary of Proposition 3.3 is that the extremal function $f_{q,r}(N)$ that we are studying is sub-multiplicative.
@@
As another consequence of Corollary 3.4, we can prove a “finitary” version of this limiting statement, similar to the use of the tensor product trick in the study of Sidorenko’s conjecture. It shows that whenever we prove a lower bound of the form $f_{q,r}(N)\geqslant N^{\alpha-o(1)}$ for some $\alpha$ and some $o(1)$ term tending to zero as $N\to\infty$, we can automatically deduce the stronger bound $f_{q,r}(N)\geqslant N^{\alpha}$ for all $N$.
@@
$\Pi(K)\coloneqq\prod_{i=1}^{q}\ell_{i}(K),$
where $\ell_{i}(K)$ denotes the length of the longest $i$-avoiding monotone path in $K$. A priori, working with this product may appear to be wasteful, since perhaps the longest color-avoiding path is much longer than the geometric mean of all of them. As an immediate consequence of Proposition 3.3, we can show that this is not the case, and that the two notions are equivalent for large $N$.
@@
For a $q$-edge-colored tournament $T$, we denote by $g_{q,r}(T)$ the maximum length of a directed path in $T$ that receives at most $r$ colors. We then let $g_{q,r}(N)$ denote the minimum of $g_{q,r}(T)$ over all $N$-vertex tournaments $T$. Note that we have the trivial bound $g_{q,r}(N)\leqslant f_{q,r}(N)$, since $f_{q,r}(N)$ precisely corresponds to restricting this minimization to transitive tournaments. The following lemma records some simple relations between these functions for different choices of the parameters $r$ and $q$. These bounds can be seen as analogues of the simple relations (2) for the vector functions $F_{q,r}$ and $G_{q,r}$.
@@
Although these bounds are very simple, they are sufficient to determine $f_{q,r}$ and $g_{q,r}$ in certain special cases. For example, recall that the Erdős–Szekeres lemma implies that $f_{q,1}(N)\geqslant N^{1/q}$, and the Gallai–Hasse–Roy–Vitaver theorem implies that $g_{q,1}(N)\geqslant N^{1/q}$; both bounds are tight if $N$ is a power of $q$. Using Lemma 3.8(2), we conclude that for all $d\geqslant 1$,
@@
3.3. Reductions: proof of Theorems 1.2 and 1.5
We now show how the various reductions proved in the previous subsection immediately imply Theorems 1.2 and 1.5, assuming Theorem 1.6. We begin with Theorem 1.5.
Proof of Theorem 1.5, assuming Theorem 1.6.
Theorem 1.6 states that every $q$-edge-colored NN-vertex tournament contains a color-avoiding directed path of length at least $c_{q}N^{\alpha}$, where $\alpha=1-O(1/\sqrt{\log q})$. In particular, applying this to transitive tournaments, we find that $f_{q,q-1}(N)\geqslant c_{q}N^{\alpha}=N^{\alpha-\varepsilon(N)}$, where $\varepsilon(N)=\log\frac{1}{c_{q}}/{\log N}$. Since $\lim_{N\to\infty}\varepsilon(N)=0$, we may apply Proposition 3.6 to conclude that, in fact, $f_{q,q-1}(N)\geqslant N^{\alpha}=N^{1-O(1/\sqrt{\log q})}$ for all $N$. This is precisely the statement of Theorem 1.5. ∎
Similarly, Theorem 1.2 follows immediately, as it is equivalent to Theorem 1.5.
Proof of Theorem 1.2, assuming Theorem 1.6.
We have already seen how Theorem 1.6 implies Theorem 1.5, hence we may assume that $f_{q,q-1}(N)\geqslant N^{\alpha}$ for all $N$, where $\alpha=1-O(1/\sqrt{\log q})$. By Proposition 3.1, this is the same as saying that $F_{q,q-1}(n)\leqslant n^{1/\alpha}$. But $1/\alpha=1+O(1/\sqrt{\log q})$, which proves Theorem 1.2. ∎
@@
(6) $\gamma=\frac{1}{2q\cdot 2^{\sqrt{\log q}}}.$
@@
(7) $p=\frac{24q}{2^{\sqrt{\log q}}},\qquad s=2\gamma N=\frac{N}{q\cdot 2^{\sqrt{\log q}}},\qquad\delta=\frac{\gamma}{8}=\frac{1}{16q\cdot 2^{\sqrt{\log q}}}.$
@@
(8) $\Pi(T)\geqslant 2^{\delta q}\cdot c_{q}\cdot N^{q-Cq/\sqrt{\log q}},$
@@
As $G$ is an $N$-vertex ordered graph, we henceforth identify the vertices of $G$ with the interval $[N]$, so that the ordering $\prec$ of $G$ agrees with the standard ordering of $N$. We alternate between these two notations as convenient. We begin by partitioning $[N]=V(G)=V(T)$ into four intervals $A,B,C,D$, of lengths $\frac{N}{2}-4s,4s,4s,\frac{N}{2}-4s$, respectively, where we recall that $s$ was defined in (7). That is, we have
@@
$\frac{\Pi(T)}{c_{q}N^{q-Cq/\sqrt{\log q}}}$ $\geqslant\frac{2^{q-2p-128}\Pi_{q}(N^{\prime})}{c_{q}N^{q-Cq/\sqrt{\log q}}}$
$\geqslant 2^{q-2p-128}\cdot\frac{(N^{\prime})^{q-Cq/\sqrt{\log q}}}{N^{q-Cq/\sqrt{\log q}}}$
@@
Let $i$ be a short color, and let $v\in X_{i}$. There is a set $E_{i}(v)\subseteq A\cup B$ with $\lvert E_{i}(v)\rvert<2s$ such that the following holds for all $w\in(A\cup B)\setminus E_{i}(v)$. Vertex $v$ is adjacent to $w$ in $G$, and if $v\to w$ in $T$, then the color of the edge $vw$ is $i$.
@@ 5. Concluding remarks
Another natural question left open by our work is to extend our results to the more general settings of $r$-comparable sets in $n]^{q}$, and of directed paths in $q$-edge-colored tournaments which receive at most $r$ colors, for more general parameters $1\leqslant r<q$. For example, Theorem 1.7 allows us to find long paths that receive only $3$ colors whenever we are working with a tournament that is far from transitive; however, most of our other arguments are quite specialized to the color-avoiding setting of r=q1r=q-1, and there appear to be serious difficulties to extending them to the most general setting. On the other hand, the simple reductions recorded in Lemma 3.8 do allow us to obtain some non-trivial results in this more general setting as a consequence of Theorem 1.6. For example, we can prove the following result, which gives a good estimate when $q/r$ is close to $1$.

### v2

@@ 3. Simple proofs
The first two authors [23, Lemma 1.3] proved the following quantitative variant of the triangle removal lemma in tournaments: There is a constant $c>0$ such that if a tournament on $N$ vertices is $\delta$-far from transitive, then it contains at least $c\delta^{2}N^{3}$ cyclic triangles. For each cyclic triangle, arbitrarily label the edges cyclically as $1,2,3$. Since there are $q$ colors, there are $q^{3}$ possible colorings of the edges of a cyclic triangle with the edges labeled as $1,2,3$. By the pigeonhole principle, there are at least $c\delta^{2}N^{3}/q^{3}$ cyclic triangles whose labeled edges have the same color pattern (call this pattern $P$). Greedily delete edges from the tournament that are in fewer than $2c\delta^{2}N/q^{3}$ such cyclic triangles of color pattern $P$. As there are fewer than $N^{2}/2$ edges, we still have more than $c\delta^{2}N^{3}/q^{3}-(N^{2}/2)(2c\delta^{2}N/q^{3})=0$ cyclic triangles of color pattern $P$ remaining after we are done with this process. Each edge that remains (and there is at least one since there is at least one cyclic triangle remaining) extends to at least $2c\delta^{2}N/q^{3}$ cyclic triangles of color pattern $P$. Starting with the vertices $v_{1},v_{2},v_{3}$ of one remaining cyclic triangle of color pattern $P$ with $v_{1}\rightarrow v_{2}$, we can greedily build the desired directed path (and furthermore the square of a path with the desired properties) by adding, after $v_{i-1},v_{i}$, a vertex $v_{i+1}$ not already on the path such that $v_{i-1},v_{i},v_{i+1}$ forms a cyclic triangle of color pattern $P$ in what remains. We can necessarily find such a vertex $v_{i+1}$ as long as $2c\delta^{2}N/q^{3}>i-1$ since each edge on the path extends to more than $2c\delta^{2}N/q^{3}$ cyclic triangles of color pattern $P$ in what remains. Hence, the path length we get in the end is at least $2c\delta^{2}N/q^{3}$. ∎
@@ 3.2. General reductions and bounds
One useful observation is that monotone paths behave very cleanly under lexicographic product. Recall that an ordered graph is a graph equipped with a linear order on its vertex set. Given two ordered graphs $G_{1},G_{2}$, we define their lexicographic product $G_{1}\otimes G_{2}$ by first duplicating each vertex of $G_{2}$ to $\lvert G_{1}\rvert$ vertices, and replacing each edge of $G_{2}$ by a complete bipartite graph between these blowup sets. We then insert a copy of $G_{1}$ into each of these blowup sets, while otherwise maintaining the order. That is, each original vertex of $G_{2}$ now corresponds to an interval of length $\lvert G_{1}\rvert$ in the new ordering.
@@
One consequence of Proposition 3.2 is the upper bound $g_{q,r}(N)\leqslant f_{q,r}(N)\leqslant N^{r/q}$ for all $N$ which are a power of $q$. Indeed, let $N=m^{q}$, and let $K_{1},\dots,K_{q}$ be ordered cliques on $m$ vertices, where all edges in $K_{i}$ are given color $i$. By Proposition 3.2, for every $S\subseteq[q]$, the length of the longest $S$-colored monotone path in $K=K_{1}\otimes\dots\otimes K_{q}$ is precisely $m^{\lvert S\rvert}$; in particular, the longest monotone path in $K$ receiving at most $r$ colors has length exactly $m^{r}=(m^{q})^{r/q}$, proving the bound $f_{q,r}(N)\leqslant N^{r/q}$ for $N=m^{q}$.
Another immediate corollary of Proposition 3.2 is that the function $f_{q,r}(N)$ is sub-multiplicative.
@@
As another consequence of Corollary 3.3, we can prove a “finitary” version of this limiting statement, similar to the use of the tensor product trick in the study of Sidorenko’s conjecture (e.g. [2, 16]). It shows that whenever we prove a lower bound of the form $f_{q,r}(N)\geqslant N^{\alpha-o(1)}$ for some $\alpha$ and some $o(1)$ term tending to zero as $N\to\infty$, we can automatically deduce the stronger bound $f_{q,r}(N)\geqslant N^{\alpha}$ for all $N$.
@@
$\Pi(T)\coloneqq\prod_{i=1}^{q}\ell_{i}(T),$
where $\ell_{i}(T)$ denotes the length of the longest $i$-avoiding directed path in $T$. A priori, working with this product may appear to be wasteful, since perhaps the longest color-avoiding path is much longer than the geometric mean of all of them. As an immediate consequence of Proposition 3.2, we can show that in the setting of transitive tournaments, this is not the case, and that the two notions are equivalent for large $N$.
@@
Finally, the following lemma records some simple relations between the quantities $f_{q,r}(N)$ and $g_{q,r}(N)$ for different choices of the parameters $r$ and $q$.
@@
Although these bounds are very simple, they are sufficient to determine $f_{q,r}$ and $g_{q,r}$ in certain special cases. For example, recall that the Chvátal–Gyárfás–Lehel theorem states that $f_{q,1}(N)\geqslant g_{q,1}(N)\geqslant N^{1/q}$; both bounds are tight if $N$ is a power of $q$. Using Lemma 3.7(2), we conclude that for all $d\geqslant 1$,
@@
$\gamma=\frac{1}{2q\cdot 2^{\sqrt{\log q}}}.$ (5)
@@
$p=\frac{24q}{2^{\sqrt{\log q}}},\qquad s=2\gamma N=\frac{N}{q\cdot 2^{\sqrt{\log q}}},\qquad\delta=\frac{\gamma}{8}=\frac{1}{16q\cdot 2^{\sqrt{\log q}}}.$ (6)
@@
$\Pi(T)\geqslant 2^{\delta q}\cdot c_{q}\cdot N^{q-Cq/\sqrt{\log q}},$ (7)
@@
As $G$ is an $N$-vertex ordered graph, we henceforth identify the vertices of $G$ with the interval $[N]$, so that the ordering $\prec$ of $G$ agrees with the standard ordering of $[N]$. We alternate between these two notations as convenient. We begin by partitioning $[N]=V(G)=V(T)$ into four intervals $A,B,C,D$, of lengths $\frac{N}{2}-4s,4s,4s,\frac{N}{2}-4s$, respectively, where we recall that $s$ was defined in (6). That is, we have
@@
$\frac{\Pi(T)}{c_{q}N^{q-Cq/\sqrt{\log q}}}$ $\geqslant\frac{2^{q-2p-128}\Pi_{q}(N^{\prime})}{c_{q}N^{q-Cq/\sqrt{\log q}}}\geqslant 2^{q-2p-128}\cdot\frac{(N^{\prime})^{q-Cq/\sqrt{\log q}}}{N^{q-Cq/\sqrt{\log q}}}$
@@
Let $i$ be a short color, and let $v\in X_{i}$. There is a set $E_{i}(v)\subseteq A\cup B$ with $\lvert E_{i}(v)\rvert<2s$ such that the following holds for all $w\in(A\cup B)\setminus E_{i}(v)$. The vertices $v$ and $w$ are adjacent in $G$, and if $v\to w$ in $T$, then the color of the edge $vw$ is $i$.
@@ 5. Concluding remarks
Another natural question left open by our work is to extend our results to the more general setting of directed paths in $q$-edge-colored tournaments which receive at most $r$ colors, for more general parameters $1\leqslant r<q$. For example, Theorem 1.2 allows us to find long paths that receive only $3$ colors whenever we are working with a tournament that is far from transitive; however, many of our other arguments seem specialized to the color-avoiding setting of r=q1r=q-1, and there appear to be difficulties in extending them to the most general setting. On the other hand, the simple reductions recorded in Lemma 3.7 do allow us to obtain some non-trivial results in this more general setting as a consequence of Theorem 1.1. For example, we can prove the following result, which gives a good estimate when $q/r$ is close to $1$.
@@
We stress that these questions are also interesting in the transitive case. In particular, while Proposition 1.6 provides good bounds on $F_{q,r}(n)$, once $r\neq q-1$, this problem is no longer equivalent to the study of $f_{q,r}(N)$, but rather to a set-colored Ramsey question. Thus, determining the behavior of $f_{q,r}(N)$ remains an interesting open problem.
Acknowledgments
We would like to thank Cosmin Pohoata for bringing the work of Sawin and Tao to our attention, and for helpful discussions on this topic.